### v1

# Extending wavelet regularity beyond Gevrey classes

Abstract
We construct a smooth orthonormal wavelet $\psi$ such that both $\psi$ and its Fourier transform $\widehat{\psi}$ belong to the extended Gevrey class $\mathcal{E}_{\sigma}(\mathbb{R})$ for $\sigma>1$, providing an example that lies beyond all classical Gevrey classes. Our approach uses the idea of invariant cycles to extend the initial Lemarié-Meyer support of the low-pass filter $m_{0}$ from $\frac{2\pi}{3},\frac{2\pi}{3}]$ to $\frac{4\pi}{5},\frac{4\pi}{5}]$. This extension allows us to control the decay rate of $m_{0}$ near $\frac{2\pi}{3}$, which yields global decay estimates for $\psi$ and $\hat{\psi}$. In addition, the decay rates are described using special functions involving the Lambert W function, which plays an important role in our construction.
1 Introduction
One of the fundamental concepts in wavelet theory is the notion of a multiresolution analysis (MRA in short), introduced by Y. Meyer [17] and later developed by S. Mallat [16]. MRA is based on a hierarchical family of closed subspaces of $L^{2}(\mathbb{R})$ and involves a scaling function $\varphi$, which plays a central role in wavelet construction. Under certain technical assumptions, $\hat{\varphi}$ can be represented as an infinite product constructed from a $2\pi$-periodic function, known as the low-pass filter. More precisely, if $\varphi$ is a scaling function and $m_{0}$ the associated low-pass filter, then
$\hat{\varphi}(\xi)=\prod_{j=1}^{\infty}m_{0}(2^{-j}\xi),\qquad\xi\in\mathbb{R},$ (1)
and the Fourier transform of the corresponding MRA wavelet $\psi$ is given by
$\hat{\psi}(\xi)=e^{i\xi/2}\,\overline{m_{0}\!\left(\frac{\xi}{2}+\pi\right)}\,\hat{\varphi}\!\left(\frac{\xi}{2}\right),\qquad\xi\in\mathbb{R}.$ (2)
With this approach, it is therefore sufficient to construct a low-pass filter with the desired regularity properties, which are inherited by $\hat{\varphi}$ and $\hat{\psi}$ through (1) and (2). The regularity of $\varphi$ and $\psi$ is then measured via Paley–Wiener type theorems, which relate the smoothness of a function to the decay of its Fourier transform. For a more detailed exposition of the general theory of wavelets and their applications, we refer to [5, 9].
We will use invariant cycles $\left\{-\frac{2\pi}{3},\frac{2\pi}{3}\right\}$ and $\left\{-\frac{2\pi}{5},\frac{2\pi}{5},-\frac{4\pi}{5},\frac{4\pi}{5}\right\}$ to extend the support of the low-pass filter $m_{0}$ on the interval $\pi,\pi)$. For this purpose, let us recall the definition (see [10]).
Definition 1.1.
Let $\rho:[-\pi,\pi)\to[-\pi,\pi):\xi\mapsto 2\xi\ (mod\,2\pi$). For $\xi\in[-\pi,\pi)$, if there exists a number $l\in\mathbb{N}$ such that $\rho^{l}(\xi)=\xi$, we call the corresponding orbit $\mathcal{O}(\xi)=\{\xi,\rho(\xi),\dots,\rho^{l-1}(\xi)\}$ an invariant cycle of length $l$, generated by $\xi$.
The notion of invariant cycles was introduced by A. Cohen to characterize orthonormality conditions for wavelets (see [5]). In [10], it is shown that for a band-limited MRA wavelet, $\prod_{\xi\in\mathcal{O}(\xi)}m_{0}(\xi)=0$ for every invariant cycle $\mathcal{O}(\xi)$. The authors also constructed a smooth low-pass filter with $m_{0}(\pm 2\pi/3)\neq 0$, yielding a non-band-limited wavelet regular only up to finite order. In contrast, Lemarié–Meyer wavelets use smooth low-pass filter that vanish in a neighborhood of $\pm 2\pi/3$, making the corresponding scaling function $\varphi$ and wavelet $\psi$ band-limited and thus analytic in time.
It is well known that a smooth orthonormal wavelet cannot have exponential decay (see [6]). Therefore, constructing a smooth wavelet close to the “critical case”—that is, one which is slightly more regular than $C^{\infty}$, or whose decay is slightly slower than exponential—appears to be a challenging task. An example of such a wavelet can be found in [7], where the authors employ Gevrey regularity and use the translates of the Gevrey function $e^{-|\xi|^{-\frac{1}{s-1}}}$, $s>1$, to control the decay rate of the low-pass filter near $\frac{2\pi}{3}$.
We apply ideas from the cited work to the function
$\begin{split}&f_{\rho,\sigma}(x)=e^{-\rho\,g_{\sigma}\left(1/x\right)},\qquad\rho>0,\;\sigma>1,\\ &\text{where}\\ &g_{\sigma}(x)=\omega_{\sigma}\!\left(\ln\left(1+|x|\right)\right),\qquad x\in\mathbb{R},\\ &\omega_{\sigma}(x)=x^{\frac{\sigma}{\sigma-1}}\big/W^{\frac{1}{\sigma-1}}(x),\qquad x\geq 0,\;\sigma>1,\qquad\omega_{\sigma}(0)=0,\end{split}$ (3)
and $W$ denotes the principal branch of the Lambert $W$ function. This approach yields a new non-band-limited wavelet whose decay rate is faster than polynomial but slower than subexponential. Moreover, its regularity is captured by the extended Gevrey classes.
Extended Gevrey classes of locally smooth functions were introduced in [20]. They contain the union of all Gevrey classes and therefore describe regularity intermediate between Gevrey and $C^{\infty}$. The growth of the derivatives of their elements is controlled by the sequence $\{M^{\tau,\sigma}_{p}\}_{p\in\mathbb{N}_{0}}$, where
$M^{\tau,\sigma}_{p}=p^{\tau p^{\sigma}},\quad p\in\mathbb{N},\qquad M_{0}^{\tau,\sigma}=1,\qquad\tau>0,\ \sigma>1.$ (4)
Since the sequences in (4) do not satisfy Komatsu’s condition
$(M.2)\quad(\exists\,C>0)\ M_{p+q}\leq C^{p+q+1}M_{p}M_{q},\quad p,q\in\mathbb{N},$
for any choice of $\tau>0$ and $\sigma>1$, they cannot be treated within the classical framework of ultradifferentiable function theory. Nevertheless, it has recently been shown that extended Gevrey classes form an important example of weight matrix classes, where regularity is governed by a family of sequences rather than a fixed one (see [11, 23]). Applications of the theory can be found in [3, 12, 13, 2], where extended Gevrey classes are often referred as PTT-spaces (Pilipović, Teofanov, Tomić spaces).
The main purpose of this paper is to establish new regularity properties of wavelets that go beyond the classical theory of ultradifferentiable functions. We refine the result of [24] on band-limited wavelets in extended Gevrey settings by controlling the decay and smoothness of the scaling function in the Fourier domain. At the same time, we obtain a wavelet that is less regular than the one constructed in [7, Theorem 4.4]; equivalently, its regularity lies closer to $C^{\infty}$ (see Lemma 3.6).
Paper is organized as follows: In Subsection 1.1 we fix the notation and discuss the main properties of the Lambert $W$ function. In Section 2, we introduce the extended Gevrey classes $\mathcal{E}_{\sigma}(\mathbb{R})$, discuss their defining sequences $M^{\tau,\sigma}_{p}$ and corresponding associated functions. In Section 2.1 we prove that functions $f_{\rho,\sigma}$ given in (3) belong to $\mathcal{E}_{\sigma}(\mathbb{R})$. Finally, in Section 3 we construct the desired wavelet from the low-pass filter $m_{0}$, by using $f_{\rho,\sigma}$, and prove Theorem 1.2 as our main result:
Theorem 1.2.
Let $\sigma>1$. For $\eta>1$ set $\sigma_{\eta}=\frac{\sigma+\eta(\sigma-1)}{1+\eta(\sigma-1)}\in(1,\sigma_{0})$ where $\sigma_{0}=\max\{2,\sigma\}$. Then there exists an orthonormal wavelet $\psi$ such that
$\psi\in\mathcal{E}_{\sigma}(\mathbb{R})\backslash\bigcup_{1<\sigma^{\prime}<\sigma_{\eta}}\mathcal{E}_{\sigma^{\prime}}(\mathbb{R})\quad{\rm and}\quad\displaystyle\hat{\psi}\in\mathcal{E}_{\sigma}(\mathbb{R})\backslash\bigcup_{1<\sigma^{\prime}<\sigma}\mathcal{E}_{\sigma^{\prime}}(\mathbb{R}).$
The Appendix contains the proof of Lemma 2.17 and illustrations of the wavelet constructed in Section 3.
1.1 Preliminaries
Throughout the paper we use the following notation: $\mathbb{N}$, $\mathbb{N}_{0}$, $\mathbb{R}_{+}$, $\mathbb{R}$ and $\mathbb{C}$ denote the sets of natural numbers, non-negative integers, positive real numbers, real numbers, and complex numbers, respectively. We write $f\asymp g$ to denote that the two functions are asymptotically equivalent, meaning that $f=O(g)$ and $g=O(f)$ as $x\to\infty$ (here $f=O(g)$, $x\to\infty$, means that $f(x)\leq L\,(g(x)+1)$ for some $L\geq 1$ and all $x\geq 0$). By $f\prec g$ we mean that $f=o(g)$ as $x\to\infty$, that is, $f(x)/g(x)\to 0$ as $x\to\infty$.
With ${\rm coz}\,f=\{x_{0}\in\mathbb{R}\,|\,f(x_{0})\not=0\}$ we denote a cozero set (complement of a zero set) of the continuous function $f$. Then the support of $f$ is ${\rm supp}\,f=\overline{{\rm\operatorname{coz}}\,f}$, where $\overline{X}$ denotes the closure of the set $X$. The interior of the set $X$ is denoted by ${\rm int}\,X$. We write $\mathbb{K}\subset\subset U$ when $\mathbb{K}$ is an compact subset of an open set $U$.
Let $\nu$ be a non-negative, continuous, increasing and even function on $\mathbb{R}$, with $\nu(0)=0$. Recall (see [1]), $\nu$ is of Braun-Meise-Taylor type (BMT in short) if the following conditions hold:
($\alpha$)   $\nu(2x)=O(\nu(x)),\quad x\to\infty,$
($\beta$)   $\nu(x)=O(x),\quad x\to\infty,$
($\gamma$)   $\ln x=o(\nu(x)),\quad x\to\infty,\;\;$
($\delta$)   $\vartheta(x)=\nu(e^{x})\quad$ is convex.
Example 1.3.
Some classical examples of BMT functions are
$\nu(x)=|x|^{s},\,0<s\leq 1,\quad\nu(x)=\ln^{s}_{+}|x|,\,s>1,\quad\nu(x)=\frac{|x|}{\ln^{s-1}(e+|x|)},\,s>1,$
for $x\in\mathbb{R}$ where $\ln_{+}{|x|}=\max\{0,\ln|x|\}$. Moreover, it was shown in [23] that $\omega_{\sigma}(\ln_{+}|x|)$, $x\in\mathbb{R}$, where $\omega_{\sigma}$ is given in (3), is asymptotically equivalent to a BMT function. Clearly, the same conclusion holds for $g_{\sigma}(x)=\omega_{\sigma}(\ln(1+|x|))$.
We recall some of the basic properties of the Lambert $W$ function. It is defined as the multivalued inverse of the function $z\mapsto ze^{z}$ for $z\in\mathbb{C}$. We will denote its principal branch by $W(z)$, $z\in\mathbb{C}\setminus(-\infty,-e^{-1}]$ (instead of common $W_{0}(z)$). Lambert $W$ function splits the complex plane into infinitely many regions. The boundary curve of its principal branch is given by:
$\{(-x\,\operatorname{ctg}x,x)\in\mathbb{R}^{2}\,|\,-\pi<x<\pi\}.$
Here we list some additional properties of $W$ that we will use in the sequel:
  • $(W1)\quad$
    $W(-e^{-1})=-1$, $W(0)=0$, $W(e)=1$ and $W(x)$ is continuous, increasing and concave on $(-e^{-1},\infty)$.
  • $(W2)\quad$
    $z=W(z)e^{W(z)}$, $z\in\mathbb{C}\setminus(-\infty,-e^{-1}]$.
  • $(W3)\quad$
    $\ln x-\ln(\ln x)\leq W(x)\leq\ln x-\frac{1}{2}\ln(\ln x),\quad x\geq e.$
  • $(W4)\quad$
    Derivatives of Lambert $W$ function are given by
    $W^{(n)}(x)=\frac{W^{n}(x)\,p_{n}(W(x))}{x^{n}(1+W(x))^{2n-1}},\quad x>0,\quad n\in\mathbb{N},$
    where $p_{n}$ is polynomial satisfying
    $p_{n+1}(x)=(1+x)p^{\prime}_{n}(x)-(nx+3n-1)p_{n}(x),\quad p_{1}(x)=1.$
Note that $(W1)$ implies that $W(x)>0$ when $x>0$. From $(W3)$ it follows
$W(x)\sim\ln x,\quad x\to\infty,$ (5)
and hence $W(Cx)\sim W(x),$ $x\to\infty$, for any $C>0$. Moreover, by $(W1)$ and $(W2)$ we have
$W(x)\sim x,\quad x\to 0,$ (6)
and $(W4)$ implies that
$W^{\prime}(x)=\frac{W(x)}{x(1+W(x))},\quad x>0.$ (7)
For more details concerning the Lambert $W$ function we refer to [4, 18].
We end this preliminary section with the following lemma that describes the behavior of $e^{W(x)}$.
Lemma 1.4.
If $W$ is the principal branch of Lambert $W$ function then
$\ln x\prec e^{W(x)}\prec x,\quad x\to\infty.$
Proof.
By using L’Hôpital’s rule, (5) and (7), we have
$\lim_{x\to+\infty}\frac{\ln x}{e^{W(x)}}=\lim_{x\to+\infty}\frac{\frac{1}{x}}{e^{W(x)}\frac{W(x)}{x(1+W(x))}}=\lim_{x\to+\infty}\frac{1}{e^{W(x)}}\frac{1+W(x)}{W(x)}=0.$
For the second part, we use $(W2)$ and (5) to obtain
$\lim_{x\to+\infty}\frac{e^{W(x)}}{x}=\lim_{x\to+\infty}\frac{e^{W(x)}}{W(x)e^{W(x)}}=\lim_{x\to+\infty}\frac{1}{W(x)}=0.\qed$
2 Regularity classes
We begin with a Lemma from [20] that summarizes the main properties of the sequences $M^{\tau,\sigma}_{p}$, $p\in\mathbb{N}_{0},$ given in (4).
Lemma 2.1.
Let $\tau>0$, $\sigma>1$, $M_{0}^{\tau,\sigma}=1$, and $M_{p}^{\tau,\sigma}=p^{\tau p^{\sigma}}$, $p\in\mathbb{N}$. Then there exists constant $C>1$ such that:
  • $(M.1)$
    $(M_{p}^{\tau,\sigma})^{2}\leq M_{p-1}^{\tau,\sigma}M_{p+1}^{\tau,\sigma},\quad$ $p\in\mathbb{N},$
  • $\widetilde{(M.2)}$
    $M_{p+q}^{\tau,\sigma}\leq C^{p^{\sigma}+q^{\sigma}}M_{p}^{\tau 2^{\sigma-1},\sigma}M_{q}^{\tau 2^{\sigma-1},\sigma},\quad$ $p,q\in\mathbb{N}_{0},\quad$
  • $\widetilde{(M.2)^{\prime}}$
    $M_{p+1}^{\tau,\sigma}\leq C^{p^{\sigma}}M_{p}^{\tau,\sigma},\quad$ $p\in\mathbb{N}_{0},\quad$
  • $(M.3)^{\prime}$
    $\sum\limits_{p=1}^{\infty}\frac{M_{p-1}^{\tau,\sigma}}{M_{p}^{\tau,\sigma}}<\infty.$
In addition, if $\sigma_{2}>\sigma_{1}>1$ and $\tau_{0}>0$ then for every $h,\tau>0$ there exists $C>0$ such that
$h^{p^{\sigma_{1}}}M^{\tau_{0},\sigma_{1}}_{p}\leq CM_{p}^{\tau,\sigma_{2}}.$ (8)
Remark 2.2.
Let us briefly comment the case $\sigma=1$. Note that $M^{\tau,1}_{p}=p^{\tau p}$, $p\in\mathbb{N}$, for $\tau>0$ are Gevrey sequences. Then the conditions $\widetilde{(M.2)^{\prime}}$ and $\widetilde{(M.2)}$ reduce to the classical Komatsu conditions $(M.2)^{\prime}$} and (M.2){(M.2)}, respectively. Moreover, the Gevrey sequences satisfy non-quasianalyticity condition $(M.3)^{\prime}$} if and only if $\tau>1$.
Now we can define the extended Gevrey classes $\mathcal{E}_{\tau,\sigma}(\mathbb{R})$ for $\tau>0$ and $\sigma>1$.
Definition 2.3.
Let $\tau>0$ and $\sigma>1$ and $M^{\tau,\sigma}_{p}=p^{\tau p^{\sigma}}$ for $p\in\mathbb{N}$, $M^{\tau,\sigma}_{0}=1$. A smooth function $\phi$ belongs to $\mathcal{E}_{\tau,\sigma}(\mathbb{R})$ if
$(\forall\,\mathbb{K}\subset\subset\mathbb{R})(\exists C>0)\quad\sup_{x\in\mathbb{K}}|\phi^{(p)}(x)|\leq C^{p^{\sigma}+1}M^{\tau,\sigma}_{p},\,\,p\in\mathbb{N}_{0}.$ (9)
We denote
$\displaystyle\mathcal{E}_{\sigma}(\mathbb{R})=\bigcup_{\tau>0}\mathcal{E}_{\tau,\sigma}(\mathbb{R}).$
Remark 2.4.
In the definition of $\mathcal{E}_{\tau,\sigma}(\mathbb{R})$ the space $\mathbb{R}$ may be replaced by any open set $U\subseteq\mathbb{R}$. This produces the local classes $\mathcal{E}_{\tau,\sigma}(U)$; their elements satisfy derivative estimates (9) on every compact set $\mathbb{K}\subset\subset U$.
Again for $\sigma=1$ we obtain some of the well-known classes. For instance, $\mathcal{E}_{1,1}(\mathbb{R})=\mathcal{A}(\mathbb{R})$, and $\mathcal{E}_{\tau,1}(\mathbb{R})=\mathcal{G}_{\tau}(\mathbb{R})$ for $\tau>1$; these are the spaces of locally analytic functions and Gevrey functions of order $\tau$, respectively.
Extended Gevrey classes $\mathcal{E}_{\tau,\sigma}$ for $\sigma>1$ and $\tau>0$ were introduced and studied as spaces of locally smooth functions equipped with projective and inductive limit topologies. Since these topologies will not be used in this paper, we omit the details and refer the reader to [20]. Unlike the classical Gevrey classes, note that in the definition of $\mathcal{E}_{\tau,\sigma}$ the geometric factor $C^{p}$ is replaced by $C^{p^{\sigma}}$, $\sigma>1$. This modification ensures the stability of the classes under the action of differential operators.
Taking the union with respect to the parameter $\tau$, we obtain the weight-matrix class $\mathcal{E}_{\sigma}$ of ultradifferentiable functions, whose derivatives are controlled by the family of sequences $\{M^{\tau,\sigma}_{p}\}_{\tau>0,\,p\in\mathbb{N}_{0}}$ for $\sigma>1$. These classes can also be treated within the framework of BMT (Braun–Meise–Taylor) theory (see [23, 11]).
We summarize several basic properties of the extended Gevrey classes in the following Proposition (see [20, 21, 22, 23, 24, 25]); the proof relies on the properties of the sequences $M^{\tau,\sigma}_{p}$ stated in Lemma 2.5.
Proposition 2.5.
Let $\tau>0$ and $\sigma>1$.
  • $i)$
    For $\sigma_{2}>\sigma_{1}>1$ we have
    $\mathcal{A}(\mathbb{R})\subset\bigcup_{\tau>1}{\mathcal{G}}_{\tau}({\mathbb{R}})\subset\bigcap_{\tau>0}{\mathcal{E}}_{\tau,\sigma_{1}}({\mathbb{R}})\subset{\mathcal{E}}_{\tau,\sigma_{1}}({\mathbb{R}})\subset{\mathcal{E}}_{\sigma_{1}}\subset\mathcal{E}_{\sigma_{2}}\subset C^{\infty}({\mathbb{R}}).$
  • ii)ii)
    there exists a compactly supported function $\varphi\in{\mathcal{E}}_{\tau,\sigma}({\mathbb{R}})$.
  • $iii)$
    ${\mathcal{E}}_{\tau,\sigma}({\mathbb{R}})$ is closed under the pointwise multiplication.
  • $iv)$
    ${\mathcal{E}}_{\tau,\sigma}({\mathbb{R}})$ is closed under finite order derivation.
  • $v)$
    $\mathcal{E}_{\tau,\sigma}({\mathbb{R}})$ is closed under superposition. In particular, if $F(x)\in\mathcal{A}(\mathbb{R})$ and $f(x)\in\mathcal{E}_{\tau,\sigma}({\mathbb{R}})$ then $F(f(x))\in\mathcal{E}_{\tau,\sigma}({\mathbb{R}})$.
  • $vi)$
    $\mathcal{E}_{\tau,\sigma}({\mathbb{R}})$ is invariant under translations and dilatations.
Next we define a function associated to $M^{\tau,\sigma}_{p}$, $p\in\mathbb{N}_{0}$.
Definition 2.6.
Let $\tau>0$ and $\sigma>1$. Then the associated function to the sequence $M^{\tau,\sigma}_{p}$, $p\in\mathbb{N}_{0}$, is given by
$\displaystyle T_{\tau,\sigma}(x)=\sup_{p\in\mathbf{N}_{0}}\ln\frac{x^{p}}{M^{\tau,\sigma}_{p}},\quad\;\;x>0.$
Remark 2.7.
Since
$\left(M^{\tau,\sigma}_{p}\right)^{1/p}=p^{\tau p^{\sigma-1}}$
is clearly bounded below by a positive constant for any $\tau>0$ and $\sigma>1$, it follows that $T_{\tau,\sigma}(x)$ vanishes for sufficiently small $x>0$ (see [14]). Therefore we may assume $T_{\tau,\sigma}(0)=0$.
It turns out that for any fixed $\tau>0$, $T_{\tau,\sigma}(|x|)$ is asymptotically equivalent to the function $g_{\sigma}(x)=\frac{\ln^{\frac{\sigma}{\sigma-1}}(1+x)}{W^{\frac{1}{\sigma-1}}(\ln(1+x))}$ which appears in (3). In particular, following Proposition holds (see [22, 23]).
Proposition 2.8.
For any $\tau>0$ and $\sigma>1$ there exists $A_{\sigma},B_{\sigma}>0$ and $\widetilde{A}_{\tau,\sigma},\widetilde{B}_{\tau,\sigma}\in\mathbb{R}$ such that for $x>0$
$B_{\sigma}\tau^{-\frac{1}{\sigma-1}}\frac{\ln^{\frac{\sigma}{\sigma-1}}(1+x)}{W^{\frac{1}{\sigma-1}}(\ln(1+x))}+\widetilde{B}_{\tau,\sigma}\leq T_{\tau,\sigma}(x)\leq A_{\sigma}\tau^{-\frac{1}{\sigma-1}}\frac{\ln^{\frac{\sigma}{\sigma-1}}(1+x)}{W^{\frac{1}{\sigma-1}}(\ln(1+x))}+\widetilde{A}_{\tau,\sigma}.$
Remark 2.9.
Let us define
$h_{\tau,\sigma}(x):=e^{-T_{\tau,\sigma}(1/x)}=\inf_{p\in\mathbb{N}_{0}}M^{\tau,\sigma}_{p}x^{p},\quad x>0.$
Then Proposition 2.8 implies that for any $\rho>0$ and $\sigma>1$ there exist constants $A,B,\tau_{1},\tau_{2}>0$ such that
$Ah_{\tau_{1},\sigma}(x)\leq f_{\rho,\sigma}(x)\leq Bh_{\tau_{2},\sigma}(x),\quad x>0,$ (10)
where $f_{\rho,\sigma}(x)=e^{-\rho g_{\sigma}(1/x)}$ is given in (3).
Note that the right-hand side of (10) implies that for any given $\rho>0$ there exist $B_{1},\tau>0$ such that
$\sup_{x>0}\frac{f_{\rho,\sigma}(x)}{x^{p}}\leq B_{1}M^{\tau,\sigma}_{p},\quad p\in\mathbb{N}_{0}.$ (11)
The following definition introduces the decay rate studied in this paper.
Definition 2.10.
Let $\sigma>1$ and $g_{\sigma}$ be as in (3). A continuous function $\phi$ belongs to ${{\mathbf{\Gamma}}}_{\sigma}\left(\mathbb{R}\right)$ if
$(\exists\rho>0)\,(\exists C>0)\quad\quad|\phi(\xi)|\leq C\,e^{-\rho g_{\sigma}(\xi)},\quad\xi\in\mathbb{R}.$ (12)
We can now formulate the Paley–Wiener theorem, which relates the regularity of $\mathcal{E}_{\sigma}$ to the decay rate given by ${{\mathbf{\Gamma}}}_{\sigma}$. For further details, we refer to [22, 25].
Theorem 2.11.
Let $\sigma>1$ and ${{\mathbf{\Gamma}}}_{\sigma}$ be as in Definition 2.10.
  • a)
    Let $\varphi$ be a $C^{\infty}$ function with compact support. If $\varphi\in\mathcal{E}_{\sigma}(\mathbb{R})$ then $\widehat{\varphi}\in{{\mathbf{\Gamma}}}_{\sigma}(\mathbb{R})$.
  • b)
    If the function $\varphi$ is such that $\widehat{\varphi}\in{\mathbf{\Gamma}}_{\sigma}(\mathbb{R})$ then $\varphi\in\mathcal{E}_{\sigma}(\mathbb{R})$.
Remark 2.12.
Part $a)$ of Theorem 2.11 can be easily extended to smooth functions that are not compactly supported but have sufficient decay at infinity. In particular, if a smooth function $\varphi$ satisfies $\widehat{\varphi}\notin\mathbf{\Gamma}_{\sigma}$, then $\varphi\notin\mathcal{E}_{\sigma}$.
We conclude this section with the following technical lemma, which will be used later.
Lemma 2.13.
Let $\sigma>1$ and ${{\mathbf{\Gamma}}}_{\sigma}$ be as in Definition 2.10. Then $\phi(\xi)\in\mathbf{\Gamma}_{\sigma}(\mathbb{R})$ if and only if $(1+|\xi|)^{a}\phi(\xi)\in\mathbf{\Gamma}_{\sigma}(\mathbb{R})$ for arbitrary $a\in\mathbb{R}$.
Proof.
Let $\phi\in\mathbf{\Gamma}_{\sigma}(\mathbb{R})$ and $a\in\mathbb{R}$. Then by (12), exist $\rho>0,C>0$ such that
$(1+|\xi|)^{a}|\phi(\xi)|$ $\leq(1+|\xi|)^{a}C\exp\left\{-\rho\frac{\ln^{\frac{\sigma}{\sigma-1}}(1+|\xi|)}{W^{\frac{1}{\sigma-1}}(\ln(1+|\xi|))}\right\}$
$=C(1+|\xi|)^{a}e^{-\rho\,g_{\sigma}(\xi)}$
$\leq C\max\{1,2^{a}|\xi|^{a}\}e^{-\frac{\rho}{2}\,g_{\sigma}(\xi)}e^{-\frac{\rho}{2}\,g_{\sigma}(\xi)}$
$\leq C_{1}e^{-\frac{\rho}{2}\,g_{\sigma}(\xi)},\quad\xi\geq\frac{1}{2},$
for $C_{1}:=C\sup_{\xi\neq 0}\{\max\{1,2^{a}|\xi|^{a}\}e^{-\frac{\rho}{2}\,g_{\sigma}(\xi)}\}.$ Note that $C_{1}$ is finite, because
$\sup_{\xi\neq 0}|\xi|^{a}e^{-\frac{\rho}{2}\,g_{\sigma}(\xi)}=\sup_{t>0}\frac{f_{\frac{\rho}{2},\sigma}(t)}{t^{a}}<\infty,$
which follows easily by (11). The other implication can be proved in a similiar way. ∎
Remark 2.14.
In [12], it was noted that the sequence
$M^{\tau,\sigma}_{p}=p^{\tau p^{\sigma}},\qquad p\in\mathbb{N}_{0},$
satisfies a strong non-quasianalyticity condition for every choice of $\tau>0$ and $\sigma>1$:
$\sum_{q=p}^{\infty}\frac{M_{q}^{\tau,\sigma}}{(q+1)\,M_{q+1}^{\tau,\sigma}}\leq C\,\frac{M_{p}^{\tau,\sigma}}{M_{p+1}^{\tau,\sigma}},\qquad p\in\mathbb{N}_{0}.$
Moreover, if $a^{\tau,\sigma}_{p}=\frac{M^{\tau,\sigma}_{p+1}}{M^{\tau,\sigma}_{p}}$, they proved the stronger assertion that the gamma index of the sequence $\mathbb{M}^{\tau,\sigma}=\{M^{\tau,\sigma}_{p}\}_{p\in\mathbb{N}_{0}}$,
$\gamma(\mathbb{M}^{\tau,\sigma})=\sup\left\{\gamma>0:\;(\exists C>0)(\forall p\in\mathbb{N}_{0})(\forall q\geq p)\,\frac{a^{\tau,\sigma}_{p}}{(p+1)^{\gamma}}\leq C\,\frac{a^{\tau,\sigma}_{q}}{(q+1)^{\gamma}}\right\},$
satisfies $\gamma(\mathbb{M}^{\tau,\sigma})=\infty$.
This implies that the corresponding function classes admit (ultraholomorphic) extensions to unbounded sectors in $\mathbb{C}$ of the form
$S_{\gamma}:=\left\{z\in\mathcal{R}:|\arg(z)|<\frac{\gamma\pi}{2}\right\},$ (13)
for arbitrary $\gamma>0$, where $\mathcal{R}$ denotes the Riemann surface of the logarithm. Such extensions will be used in Section 2.1 to construct an example in $\mathcal{E}_{\sigma}(\mathbb{R})$.
2.1 An example of the extended Gevrey function
First, we discuss the complex extension of the functions defined in (3). For our purposes, it is sufficient to consider unbounded sectors $S_{\gamma}$ given in (13) with $0<\gamma\leq 2$, taking into account only the principal branch of the logarithm.
If $T$ and $S$ are two unbounded sectors, we say that $T$ is a proper subsector of $S$ if $\overline{T}\subseteq S\cup\{0\}$. By taking $\mathcal{R}$ as our initial set, we ensure that the vertex 0 is not included in T{T}. For example, note that $S_{2}=\mathbb{C}\setminus(-\infty,0]$ and $S_{1}=\{z\in\mathbb{C}\mid\mathrm{Re}(z)>0\}$; in this case, $S_{1}$ is a proper subsector of $S_{2}$.
We will use the following Lemma from [12].
Lemma 2.15.
Let $\tau>0$ and $\sigma>1$. Set $a_{\tau,\sigma}=\left(\frac{\sigma-1}{\tau\sigma}\right)^{\frac{1}{\sigma-1}}$ and $b_{\tau,\sigma}=\frac{\sigma-1}{\tau\sigma}\,e^{\frac{\sigma-1}{\sigma}}$.
  • a)
    Function $g_{\tau,\sigma}(x)=\frac{\ln^{\frac{\sigma}{\sigma-1}}(1+x)}{W^{\frac{1}{\sigma-1}}(b_{\tau,\sigma}\ln(1+x))}$ is positive and strictly increasing on $\mathbb{R}_{+}$.
  • b)
    Set
    $e_{\tau,\sigma}(z)=z\exp\left\{-a_{\tau,\sigma}W^{-\frac{1}{\sigma-1}}(b_{\tau,\sigma}\ln(1+z))\ln^{\frac{\sigma}{\sigma-1}}(1+z)\right\},\quad z\in S_{2},$
    where $\ln(\cdot)$ denotes a principal branch of $\rm Ln$$\cdot$]. Then for every proper subsector $T$ in $S_{2}$ there exists $C_{1},C_{2},K_{1},K_{2}>0$ such that
    $C_{2}e_{\tau,\sigma}(K_{2}|z|)\leq|e_{\tau,\sigma}(z)|\leq C_{1}e_{\tau,\sigma}(K_{1}|z|),\quad z\in T.$ (14)
In the next lemma we discuss properties of the functions given in (3).
Lemma 2.16.
Let $\sigma>1$.
  • a)
    Function $g_{\sigma}(x)=\omega_{\sigma}(\ln(1+x))=\frac{\ln^{\frac{\sigma}{\sigma-1}}(1+x)}{W^{\frac{1}{\sigma-1}}(\ln(1+x))}$ is positive and strictly increasing on $\mathbb{R}_{+}$. Moreover, for any $C>0$ we have
    $g_{\sigma}(Cx)\asymp g_{\sigma}(x),\quad x\to\infty.$ (15)
  • b)
    For every $a,b>0$
    $\lim_{x\to\infty}e^{ax}e^{-b\omega_{\sigma}(x)}=0,\quad\omega_{\sigma}(x)\sim x,\,\,x\to 0,\quad\lim_{x\to 0^{+}}\omega_{\sigma}(x)=0.$
    In particular, $\omega_{\sigma}(x)=\frac{x^{\frac{\sigma}{\sigma-1}}}{W^{\frac{1}{\sigma-1}}(x)}$ is continuous on $0,+\infty)$.
  • c)
    Function $f_{\rho,\sigma}(x)=\exp\{-\rho g_{\sigma}(1/x)\}$ is strictly increasing and continuous on $0,\infty)$ with $0\leq|f_{\rho,\sigma}(x)|<1$.
    For every $\varepsilon>0$
    $\int_{0}^{\varepsilon}f_{\rho,\sigma}(x)\,dx<\infty.$
    Moreover, a family $\{f_{\rho,\sigma}(x)\},\rho>0,\sigma>1$ is decreasing with respect to $\rho$, and increasing with respect to $\sigma$ for every fixed $x\in\mathbb{R}_{+}$.
  • d)
    For every proper subsector $T$ in $S_{2}$ there exist $A_{1},A_{2},\rho^{\prime},\rho^{\prime\prime}>0$ such that
    $C_{2}f_{\rho^{\prime\prime},\sigma}\left(|z|\right)\leq\left|f_{\rho_{\sigma},\sigma}(z)\right|\leq C_{1}f_{\rho^{\prime},\sigma}\left(|z|\right),\quad z\in T,$ (16)
    where $\rho_{\sigma}=e^{-1/\sigma}$ and
    $\displaystyle f_{\rho_{\sigma},\sigma}(z)=\exp\left\{-e^{-\frac{1}{\sigma}}W^{-\frac{1}{\sigma-1}}\left(\ln\left(1+\frac{1}{z}\right)\right)\ln^{\frac{\sigma}{\sigma-1}}\left(1+\frac{1}{z}\right)\right\},\quad z\in S_{2}.$
Proof.
Throughout the proof we set $\tau_{\sigma}=\frac{\sigma-1}{\sigma}\,e^{\frac{\sigma-1}{\sigma}}$.
$a)$ Note that $g_{\sigma}=g_{\tau_{\sigma},\sigma}$, where $g_{\tau,\sigma}$ is the function introduced in Lemma 2.15. Therefore, $g_{\sigma}$ is positive and strictly increasing. Moreover, [23, Theorem 3.1] implies that $g_{\sigma}(k)=\omega_{\sigma}(\ln(1+|k|))$, $k>0$, is equivalent to a BMT function (see also Example 1.3). Hence it satisfies the BMT condition $(\alpha)$, and then (15) follows by successive application of this condition.
$b)$ By the Lemma 1.4 and $(W2)$ it is obvious that
$\lim_{x\to+\infty}\exp\left\{ax-b\frac{x^{\frac{\sigma}{\sigma-1}}}{W^{\frac{1}{\sigma-1}}(x)}\right\}=\lim_{x\to+\infty}\exp\left\{x\left(a-b\left(\frac{x}{W(x)}\right)^{\frac{1}{\sigma-1}}\right)\right\}$
$=\lim_{x\to+\infty}\exp\left\{x\left(a-b\,e^{\frac{1}{\sigma-1}W(x)}\right)\right\}=0.$
Moreover, by (6)
$\lim_{x\to 0}\frac{\omega_{\sigma}(x)}{x}=\lim_{x\to 0}\frac{x^{\frac{\sigma}{\sigma-1}-1}}{W^{\frac{1}{\sigma-1}}(x)}=\lim_{x\to 0}\left(\frac{x}{W(x)}\right)^{\frac{1}{\sigma-1}}=1.$
$c)$ The fact that $f_{\rho,\sigma}$ is strictly increasing and continuous on $0,+\infty)$, follows from $a)$ and $b)$. In particular, for $\varepsilon>0$ we have
$\int_{0}^{\varepsilon}e^{-\rho\omega_{\sigma}\left(\ln\left(1+\frac{1}{x}\right)\right)}\,dx=\int_{\frac{1}{\varepsilon}}^{+\infty}e^{-\rho\omega_{\sigma}\left(\ln\left(1+s\right)\right)}\frac{1}{s^{2}}\,ds\leq\int_{\frac{1}{\varepsilon}}^{+\infty}\frac{ds}{s^{2}}<+\infty.$
The fact that $f_{\rho,\sigma}$ is a decreasing function in terms of $\rho$, it follows directly from the definition of $f_{\rho,\sigma}$ in (3). For a fixed $x\in\mathbb{R}$, we note
$F_{x}(\sigma):=g_{\sigma}(x)>0,\quad\sigma>1.$
Then we have
$\frac{d}{d\sigma}F_{x}(\sigma)=\frac{F_{x}(\sigma)}{(\sigma-1)^{2}}\ln\frac{W\left(\ln\left(1+|x|\right)\right)}{\ln\left(1+|x|\right)}<0,$
because $x\geq W(x)$ for all $x>0$. This implies that $f_{\rho,\sigma}$ is an increasing function in terms of $\sigma$.
$d)$ Note that $e_{\tau_{\sigma},\sigma}(z)=\displaystyle zf_{\rho_{\sigma},\sigma}(1/z)$ and hence (14) implies that
$C_{2}f_{{\rho_{\sigma}},\sigma}\left(K_{2}/|z|\right)\leq|f_{{\rho_{\sigma}},\sigma}\left(1/z\right)|\leq C_{2}f_{{\rho_{\sigma}},\sigma}\left(K_{1}/|z|\right),\quad z\in T.$ (17)
Since $\arg\frac{1}{z}=-\arg z,$ $z\neq 0$, note that (17) remains valid if we put $\omega=1/z\in T$. Since $f_{{\rho_{\sigma}},\sigma}\left(K|\omega|\right)=e^{-\rho_{\sigma}g_{\sigma}(1/(K|\omega|))}$, $K>0$, we can apply (15) and (16) follows. ∎
In the next section we will construct a wavelet using a particular function from $\mathcal{E}_{\sigma}(\mathbb{R})$. For $\sigma>1$ let us consider
$f_{\rho_{\sigma},\sigma}(x)=\exp\left\{-\rho_{\sigma}\,W^{-\frac{1}{\sigma-1}}\left(\ln\left(1+\frac{1}{|x|}\right)\right)\ln^{\frac{\sigma}{\sigma-1}}\left(1+\frac{1}{|x|}\right)\right\},\quad x\in\mathbb{R},$ (18)
where $\rho_{\sigma}=e^{-1/\sigma}$. First we prove that $f_{\rho_{\sigma},\sigma}$ is smooth but not analytic on $\mathbb{R}$.
The following Lemma holds.
Lemma 2.17.
For $\sigma>1$ and $\rho_{\sigma}=e^{-1/\sigma}$, let $f_{\rho_{\sigma},\sigma}$ be as in (18). Then it holds $\lim_{x\to 0}f_{\rho_{\sigma},\sigma}^{(j)}\left(x\right)=0$ for all $j\in\mathbb{N}_{0}$.
Proof of the part Lemma 2.17 is technical and therefore given in the Appendix. In the following proposition we capture the regularity of $f_{\rho_{\sigma},\sigma}$ by means of the classes $\mathcal{E}_{\sigma}(\mathbb{R})$.
Proposition 2.18.
For $\sigma>1$ and $\rho_{\sigma}=e^{-1/\sigma}$, let $f_{\rho_{\sigma},\sigma}$ be as in (18). Then it holds $f_{\rho_{\sigma},\sigma}\in\mathcal{E}_{\sigma}(\mathbb{R})$.
Proof.
Since $f_{\rho_{\sigma},\sigma}$ is even on $\mathbb{R}$ and Lemma 2.17 holds, we estimate the derivatives of $f_{\rho_{\sigma},\sigma}$ on $\mathbb{R}_{+}$. Choose subsector $T$ in $S_{2}$ (for instance $S_{1}$) such that for all $x>0$ it holds $D=\{z\in\mathbb{C}\,\,|\,\,|z-x|\leq\frac{x}{2}\}\subset T$.
Note that $|z-x|=\frac{x}{2}$ implies $\frac{x}{2}<|z|\leq\frac{3x}{2}$. Using Cauchy’s integral formula and (16) we obtain
$\left|\frac{d^{j}}{dx^{j}}f_{\rho_{\sigma},\sigma}(x)\right|$ $=\left|\frac{j!}{2\pi i}\int_{|z-x|=\frac{x}{2}}\frac{f_{\rho_{\sigma},\sigma}(z)}{(z-x)^{j+1}}\,dz\right|\leq\frac{j!}{2\pi}\int_{|z-x|=\frac{x}{2}}\frac{|f_{\rho_{\sigma},\sigma}(z)|}{|z-x|^{j+1}}\,|dz|$
$\leq\frac{j!2^{j}}{\pi x^{j+1}}\int_{|z-x|=\frac{x}{2}}c_{1}\,f_{\rho^{\prime},\sigma}\left(K|z|\right)\,|dz|$
$\leq\frac{C_{1}j!2^{j}}{\pi x^{j+1}}\int_{|z-x|=\frac{x}{2}}f_{\rho^{\prime},\sigma}\left(\frac{3Kx}{2}\right)\,|dz|$
$=\frac{C_{1}j!2^{j}}{x^{j}}\,f_{\rho^{\prime},\sigma}\left(\frac{3Kx}{2}\right),\quad x>0,\,\,j\in\mathbb{N}_{0},$ (19)
for suitable constants $C_{1},\rho^{\prime},K>0$, where we also used that $f_{\rho,\sigma}$ are increasing.
If the function $g_{\sigma}$ is given in (3), then (15) implies
$f_{\rho^{\prime},\sigma}\left(\frac{3Kx}{2}\right)=\exp\left\{-\rho^{\prime}g_{\sigma}\Big(\frac{2}{3K}\frac{1}{x}\Big)\right\}\leq L\exp\left\{-\rho^{\prime\prime}g_{\sigma}\Big(\frac{1}{x}\Big)\right\}=Lf_{\rho^{\prime\prime},\sigma}\left(x\right),$ (20)
for $x>0$ and suitable $L,\rho^{\prime\prime}>0$.
Take $\mathbb{K}\subset\subset\mathbb{R}_{+}$. Using (2.1), (20), (11), and simple inequality $j!\leq C^{j^{\sigma}}$, $C>1$, $j\in\mathbb{N}_{0}$, for suitable $C_{2},\tau>0$ we obtain
$\left|\frac{d^{j}}{dt^{j}}f_{\rho_{\sigma},\sigma}(x)\right|\leq C_{1}(2C)^{j^{\sigma}}\frac{f_{\rho^{\prime\prime},\sigma}\left(x\right)}{x^{j}}\leq C^{j^{\sigma}+1}_{2}M^{\tau,\sigma}_{j},\quad x\in\mathbb{K},\;j\in\mathbb{N}_{0},$
and this proves the assertion. ∎
We conclude this section by proving that the choice of $\rho_{\sigma}=e^{-1/\sigma}$ in (18) is not essential.
Proposition 2.19.
Let $\sigma>1$. Then for any $\rho>0$
$f_{\rho,\sigma}\in\mathcal{E}_{\sigma}(\mathbb{R})\backslash\bigcup_{1<\sigma^{\prime}<\sigma}\mathcal{E}_{\sigma^{\prime}}(\mathbb{R}),$
where $f_{\rho,\sigma}$ are given in (3).
Proof.
Take an arbitrary $\rho>0$. It is clear that Lemma 2.17 also holds for
$f_{\rho,\sigma}=(f_{\rho_{\sigma},\sigma})^{\rho/\rho_{\sigma}}.$
Moreover, since the function $z^{\alpha}$, $\alpha>0$, is analytic on $S_{2}=\mathbb{C}\setminus(-\infty,0]$, part $v)$ of Proposition 2.5 implies that $f_{\rho,\sigma}\in\mathcal{E}_{\sigma}(\mathbb{R})$, and therefore Proposition 2.18 holds as well.
Let us now show that
$f_{\rho,\sigma}\notin\bigcup_{1<\sigma^{\prime}<\sigma}\mathcal{E}_{\sigma^{\prime}}.$
Assume the opposite, that $f_{\rho,\sigma}\in\mathcal{E}_{\sigma^{\prime}}$ for some $\sigma^{\prime}<\sigma$. Then, for a small $t>0$, using (8) we obtain
$f_{\rho,\sigma}(t)=\int_{0}^{t}\frac{(t-x)^{p-1}}{(p-1)!}f^{(p)}_{\rho,\sigma}(x)\,dx\leq\frac{t^{p}}{p!}\,C^{p^{\sigma^{\prime}}+1}M^{\tau_{0},\sigma^{\prime}}_{p}\leq C^{\prime}\,t^{p}\,M^{\tau,\sigma}_{p},\qquad p\in\mathbb{N},$
for arbitrary $\tau>0$ and some constant $C^{\prime}>0$ (depending on $\tau$). This contradicts the left-hand side of (10).
Remark 2.20.
Note that Propositions 2.5 (part (i)) and 2.19 imply that, for arbitrary $\rho>0$ and $\sigma>1$, holds
$f_{\rho,\sigma}\notin\bigcup_{t>1}\mathcal{G}_{t},$
where $\mathcal{G}_{t}$ denotes the Gevrey classes.
3 Wavelet construction
In this section, we denote by $m_{0}(\xi)$ a low-pass filter of an orthonormal wavelet satisfying the following minimal requirements:
[…]
Recall that ${\rm coz}\,m_{0}$ denotes the cozero set (complement of a zero set) of $m_{0}$. Let us start with the following results from [7, 8].
Lemma 3.1.
Let $r>0$, $n\in\mathbb{N}$, and let $m_{0}$ be a low-pass filter that satisfies conditions $i)$, ii)ii) and $iii)$ in (21) such that $\operatorname{coz}m_{0}\cap[0,\pi]\subset\left[0,\left(\frac{2}{3}+r\right)\pi\right)$. If $\varphi$ is a scaling function associated with MRA, then for $r\leq\frac{2}{15}$ it holds
$\operatorname{coz}\hat{\varphi}\cap\left[2^{n}\pi,2^{n+1}\pi\right]\subset\left(\frac{2^{n+2}}{3}\pi-\left(3+(-1)^{n}\right)r\pi,\frac{2^{n+2}}{3}\pi+\left(3-(-1)^{n}\right)r\pi\right).$
In particular, if $r=\frac{2}{15}$ then $\operatorname{coz}m_{0}\cap[0,\pi]\subset\left[0,\frac{4}{5}\pi\right)$ and
$\left|\xi-\frac{2^{n+2}}{3}\pi\right|<\frac{8}{15}\pi,\quad\xi\in\operatorname{coz}\hat{\varphi}\,\cap\left[2^{n}\pi,2^{n+1}\pi\right].$ (22)
In addition, for $\xi\in\operatorname{coz}\hat{\varphi}\,\cap\left[2^{n}\pi,2^{n+1}\pi\right]$ following inequality holds
$\inf_{\xi\in A_{n}}\left|m_{0}(\xi)\right|\left|\prod_{j=1}^{n+1}m_{0}\left(\frac{2}{3}\pi-(-1)^{n-j}\frac{\xi-\frac{2^{n+2}}{3}\pi}{2^{j}}\right)\right|\leq|\hat{\varphi}(\xi)|\\ \leq\left|\prod_{j=1}^{n+1}m_{0}\left(\frac{2}{3}\pi-(-1)^{n-j}\frac{\xi-\frac{2^{n+2}}{3}\pi}{2^{j}}\right)\right|,$ (23)
where $A_{n}=\left(\frac{\pi}{3}-\frac{3+(-1)^{n}}{2^{n+2}}\pi,\frac{\pi}{3}+\frac{3-(-1)^{n}}{2^{n+2}}\pi\right)$, $n\in\mathbb{N}$.
If $\frac{2\pi}{3}\not\in{\rm int}({\rm supp}\,\,m_{0})$ then $\varphi$ is band-limited.
Remark 3.2.
Note that $\frac{2\pi}{3}\notin\operatorname{int}(\operatorname{supp}m_{0})$ suggests that $m_{0}$ vanishes near $\frac{2\pi}{3}$; in our construction, however, we control its local decay, and obtain a non–band-limited scaling function $\varphi$.
If $\operatorname{coz}m_{0}\cap[0,\pi]\subset\left[0,\frac{4}{5}\pi\right)$, note that property (iii) in (21) implies
$|m_{0}(\xi)|=1\quad\text{for }\xi\in\left[0,\frac{\pi}{5}\right].$
Moreover, $|m_{0}(\xi)|\leq 1$ for all $\xi\in\mathbb{R}$, and since $m_{0}\big(-\frac{2\pi}{3}\big)=0$, we also have $|m_{0}\left(\frac{\pi}{3}\right)|=1$.
In the following theorem, we control the local behavior of $m_{0}(\xi)$ using the functions $f_{\rho,\sigma}$ defined in (3), to obtain a global decay estimates for $\hat{\varphi}$.
Theorem 3.3.
Let $\sigma>1$ and let $m_{0}$ be a low-pass filter that satisfies conditions $i)$, ii)ii) and $iii)$ in (21) such that $\operatorname{coz}\,m_{0}\cap[0,\pi]\subset[0,\frac{4\pi}{5})$ and $\varphi$ be an associated scaling function. Moreover, let functions $f_{\rho,\sigma}$ be as in (3) and ${{\mathbf{\Gamma}}}_{\sigma}$ as in Definition 2.10.
If there exist constants $\varepsilon,C_{0},C_{1},\rho_{1},\rho_{0}>0$ such that
$\begin{split}C_{1}f_{\rho_{1},\sigma}\left(\xi-\frac{2\pi}{3}\right)\leq|m_{0}(\xi)|\leq C_{0}f_{\rho_{0},\sigma}\left(\xi-\frac{2\pi}{3}\right),\quad\xi\in\left[\frac{2\pi}{3}-\varepsilon,\frac{2\pi}{3}+\varepsilon\right],\end{split}$ (24)
then for all $\eta>1$
$\widehat{\varphi}\in{{\mathbf{\Gamma}}}_{\sigma}\backslash\bigcup_{1<\sigma^{\prime}<\sigma_{\eta}}{{\mathbf{\Gamma}}}_{\sigma^{\prime}},$
where $\sigma_{\eta}=\frac{\sigma+\eta(\sigma-1)}{1+\eta(\sigma-1)}\in(1,\sigma_{0})$ and $\sigma_{0}=\min\{2,\sigma\}$. Consequently, $\varphi\in{{\mathcal{E}}}_{\sigma}\backslash\bigcup_{1<\sigma^{\prime}<\sigma_{\eta}}{{\mathcal{E}}}_{\sigma^{\prime}}$.
Proof.
Without loss of generality, we may choose $\varepsilon<\frac{8\pi}{15}$ so that (24) holds. Moreover, let $n_{0}\in\mathbb{N}$ be such that $\frac{8\pi}{15\cdot 2^{n_{0}}}<\varepsilon$. Then for all $\xi\in\operatorname{coz}\,\widehat{\varphi}\cap[2^{n}\pi,2^{n+1}\pi]$ and $n\geq n_{0}$, the right-hand side of (23) implies
$|\widehat{\varphi}(\xi)|$ $\leq\left|\prod_{j=1}^{n+1}m_{0}\left(\frac{2\pi}{3}-(-1)^{n-j}\frac{\xi-\frac{2^{n+2}}{3}\pi}{2^{j}}\right)\right|$
$\leq\left|\prod_{j=n_{0}}^{n+1}m_{0}\left(\frac{2\pi}{3}-(-1)^{n-j}\frac{\xi-\frac{2^{n+2}}{3}\pi}{2^{j}}\right)\right|,$ (25)
where in the second inequality we used the fact that $|m_{0}(\xi)|\leq 1,\xi\in\mathbb{R}$.
Further note that for all $n_{0}\leq j\leq n+1$ and $\xi\in\operatorname{coz}\,\widehat{\varphi}\cap[2^{n}\pi,2^{n+1}\pi]$, (22) implies
$\left|(-1)^{n-j+1}\frac{\xi-\frac{2^{n+2}}{3}\pi}{2^{j}}\right|=\frac{|\xi-\frac{2^{n+2}}{3}\pi|}{2^{j}}\leq\frac{\frac{8\pi}{15\cdot 2^{n_{0}}}}{2^{j-n_{0}}}\leq\frac{\varepsilon}{2^{j-n_{0}}},\quad n_{0}\leq j\leq n+1.$ (26)
Now from (24) we have
$\left|m_{0}\left(\frac{2\pi}{3}-(-1)^{n-j}\frac{\xi-\frac{2^{n+2}}{3}\pi}{2^{j}}\right)\right|\leq C_{0}f_{\rho_{0},\sigma}\left((-1)^{n-j+1}\frac{\xi-\frac{2^{n+2}}{3}\pi}{2^{j}}\right)$
$=C_{0}f_{\rho_{0},\sigma}\left(\frac{|\xi-\frac{2^{n+2}}{3}\pi|}{2^{j}}\right)\leq C_{0}f_{\rho_{0},\sigma}\left(\frac{\varepsilon}{2^{j-n_{0}}}\right)\quad,n_{0}\leq j\leq n+1,$ (27)
for $\xi\in\operatorname{coz}\,\widehat{\varphi}\cap[2^{n}\pi,2^{n+1}\pi]$, where for the last inequality we used (26) and the fact that $f_{\rho,\sigma}$ is even, and increasing on $\mathbb{R}_{+}$ (see part $a)$ of the Lemma 2.16).
Moreover, since $f_{\rho_{0},\sigma}\left(\frac{\varepsilon}{2^{j-n_{0}}}\right)=\exp\Big\{-\rho_{0}g_{\sigma}\left(\frac{2^{j-n_{0}}}{\varepsilon}\right)\Big\}$, using again part $a)$ of the Lemma 2.16, for $\xi\in\operatorname{coz}\,\widehat{\varphi}\cap[2^{n}\pi,2^{n+1}\pi]$ we obtain
$\sum_{j=1}^{n+1}g_{\sigma}\left(\frac{2^{j-n_{0}}}{\varepsilon}\right)$ $\geq g_{\sigma}\left(\frac{2^{n+1-n_{0}}}{\varepsilon}\right)\geq g_{\sigma}\left(\frac{1}{2^{n_{0}}\varepsilon\pi}\xi\right)\geq C_{1}g_{\sigma}\left(\xi\right),$ (28)
where the last inequality follows from (15).
Finally using (3), (3), (28), and noting that $2^{n}\pi\leq|\xi|\leq 2^{n+1}\pi$ is equivalent to $n+\log_{2}\pi\leq\log_{2}|\xi|\leq n+1+\log_{2}\pi$, we obtain
$|\widehat{\varphi}(\xi)|$ $\leq\prod_{j=n_{0}}^{n+1}C_{0}f_{\rho_{0},\sigma}\left(\frac{\varepsilon}{2^{j-n_{0}}}\right)=C\prod_{j=1}^{n+1}C_{0}f_{\rho_{0},\sigma}\left(\frac{\varepsilon}{2^{j-n_{0}}}\right)$
$=CC_{0}^{n+1}\exp\left\{-\rho_{0}\sum_{j=1}^{n+1}g_{\sigma}\left(\frac{2^{j-n_{0}}}{\varepsilon}\right)\right\}$
$\leq CC_{0}^{n+1}\exp\left\{-C_{1}\rho_{0}g_{\sigma}\left(\xi\right)\right\}$
$\leq CC_{0}^{\log_{2}|\xi|+1-\log_{2}\pi}\exp\left\{-\rho^{\prime}_{0}g_{\sigma}\left(\xi\right)\right\}$
$\leq C^{\prime}|\xi|^{\nu_{0}}\exp\left\{-\rho^{\prime}_{0}g_{\sigma}\left(\xi\right)\right\},$
where $\rho^{\prime}_{0}=C_{1}\rho_{0}$, $C^{-1}=\prod_{j=1}^{n_{0}-1}C_{0}f_{\rho_{0},\sigma}\left(\frac{\varepsilon}{2^{j-n_{0}}}\right)$, $C^{\prime}=CC_{0}^{1-\log_{2}\pi}$ and $\nu_{0}=\log_{2}C_{0}$. Now Lemma 2.13 implies that $\hat{\varphi}\in{\mathbf{\Gamma}}_{\sigma}$.
It remains to show that for any $\eta>1$, $\widehat{\varphi}\notin\bigcup_{1<\sigma^{\prime}<\sigma_{\eta}}{\mathbf{\Gamma}}_{\sigma^{\prime}}$, where $\sigma_{\eta}=\frac{\sigma+\eta(\sigma-1)}{1+\eta(\sigma-1)}\in(1,\sigma_{0})$.
It is sufficient to evaluate $\hat{\varphi}$ at the points
$\xi_{n}=\frac{2^{n+2}}{3}\pi-(-1)^{n}\varepsilon\in\operatorname{coz}\,\widehat{\varphi}\cap[2^{n}\pi,2^{n+1}\pi],\quad n\in\mathbb{N},\quad\varepsilon<\frac{8\pi}{15}.$ (29)
In particular, we will prove that for arbitrary $1<\sigma^{\prime}<\sigma_{\eta}$ and $C,\rho>0$ there exists $n^{\prime}_{0}\in\mathbb{N}$ such that
$|\widehat{\varphi}(\xi_{n})|\geq C\,\exp\{-\rho\,g_{\sigma^{\prime}}(\xi_{n})\},\quad n\geq n^{\prime}_{0},$ (30)
where $\xi_{n}$ are given in (29).
Since $m_{0}\left(\frac{\pi}{3}\right)=1$ (see Remark 3.2), there exist $n_{1}\in\mathbb{N}$ such $\inf_{\xi\in A_{n}}|m_{0}(\xi)|>0$ for all $n\geq n_{1}$, where $A_{n}$ is given in Lemma 3.1.
Set $\tilde{C}=\inf_{\xi\in A_{n}}|m_{0}(\xi)|$, and note that it depends only on $n_{1}$. From Lemma 3.1, (24) and $n+\log_{2}\pi\leq|\xi|\leq n+1+\log_{2}\pi$ for $n\geq n_{1}$ we obtain
$|\widehat{\varphi}(\xi_{n})|$ $\geq\inf_{\xi\in A_{n}}|m_{0}(\xi)|\left|\prod_{j=1}^{n+1}m_{0}\left(\frac{2\pi}{3}-(-1)^{n-j}\frac{\xi_{n}-\frac{2^{n+2}}{3}\pi}{2^{j}}\right)\right|$
$=\tilde{C}\left|\prod_{j=1}^{n+1}m_{0}\left(\frac{2\pi}{3}-\frac{(-1)^{n-j}}{2^{j}}\left(\frac{2^{n+2}}{3}\pi-(-1)^{n}\varepsilon-\frac{2^{n+2}}{3}\pi\right)\right)\right|$
$=\tilde{C}\left|\prod_{j=1}^{n+1}m_{0}\left(\frac{2\pi}{3}+(-1)^{j}\frac{\varepsilon}{2^{j}}\right)\right|$
$\geq\tilde{C}C^{n+1}_{1}\prod_{j=1}^{n+1}f_{\rho_{1},\sigma}\left((-1)^{j}\frac{\varepsilon}{2^{j}}\right)$
$\geq\tilde{C_{1}}|\xi_{n}|^{\nu_{1}}\exp\left\{-\rho_{1}\sum_{j=1}^{n+1}g_{\sigma}\left(\frac{2^{j}}{\varepsilon}\right)\right\}$ (31)
for $\tilde{C}_{1}=\tilde{C}C_{1}^{-\log_{2}\pi},\nu_{1}=\log_{2}C_{1}>0$.
Take arbitrary $\eta>1$ set $\sigma_{\eta}=\frac{\sigma+\eta(\sigma-1)}{1+\eta(\sigma-1)}\in(1,\sigma_{0})$. Moreover, if we denote $a_{j}=\ln\left(1+\frac{2^{j}}{\varepsilon}\right)$, then $g_{\sigma}\left(\frac{2^{j}}{\varepsilon}\right)=\omega_{\sigma}(a_{j})$, where $\omega_{\sigma}$ is given in (3). Hence using property $(W2)$ of the Lambert $W$ function, part $a)$ of Lemma 2.16, and that $2^{n}\pi\leq\xi_{n}\leq 2^{n+1}\pi$ we obtain
$\sum_{j=1}^{n+1}g_{\sigma}\left(\frac{2^{j}}{\varepsilon}\right)=\sum_{j=1}^{n+1}\omega_{\sigma}\left(a_{j}\right)=\sum_{j=1}^{n+1}a_{j}\exp\left\{{\frac{1}{\sigma-1}W(a_{j})}\right\}$
$\leq a_{n+1}\exp\left\{\frac{1+\eta(\sigma-1)}{\sigma-1}W(a_{n+1})\right\}\cdot\sum_{j=1}^{n+1}\exp\{-\eta W(a_{j})\}$
$\leq a_{n+1}\exp\left\{\frac{1}{\sigma_{\eta}-1}W(a_{n+1})\right\}\sum_{j=1}^{\infty}\frac{W^{\eta}(a_{j})}{a^{\eta}_{j}}$
$=C_{\eta}\,g_{\sigma_{\eta}}\left(\frac{2^{n+1}}{\varepsilon}\right)\leq C_{\eta}\,g_{\sigma_{\eta}}\left(\frac{2\xi_{n}}{\pi\varepsilon}\right)\leq C^{\prime}C_{\eta}\,g_{\sigma_{\eta}}\left(\xi_{n}\right),$ (32)
for suitable $C^{\prime}>0$ where $C_{\eta}:=\sum_{j=1}^{\infty}\frac{W^{\eta}(a_{j})}{a^{\eta}_{j}}$. Let us show that $C_{\eta}<\infty$.
We need to prove that the series $\sum_{j=1}^{\infty}\frac{W^{\eta}(a_{j})}{a^{\eta}_{j}}$ is convergent. Note that
$a_{j}\sim j\ln 2-\ln\varepsilon\sim j\ln 2,\quad j\to\infty,$
and therefore by (5)
$\frac{W^{\eta}(a_{j})}{a^{\eta}_{j}}\sim\frac{\ln^{\eta}(j\ln 2)}{j^{\eta}\ln^{\eta}2}\sim C\frac{\ln^{\eta}j}{j^{\eta}},\quad j\to\infty,$
where $C=1/\ln^{\eta}2$. Clearly series $\sum\frac{\ln^{\eta}j}{j^{\eta}}$ is convergent and hence $C_{\eta}<\infty$.
Now using (3) and (3) we obtain
$|\widehat{\varphi}(\xi_{n})|\geq\tilde{C_{1}}|\xi_{n}|^{\nu_{1}}\exp\{-\rho^{\prime}_{1}\,g_{\sigma_{\eta}}(\xi_{n})\}\geq\tilde{C}_{2}\exp\{-\rho^{\prime}_{2}\,g_{\sigma_{\eta}}(\xi_{n})\},\,n\geq n_{1},$ (33)
for suitable $\tilde{C}_{2},\rho^{\prime}_{2}>0$, where we used Lemma 2.13 and $\xi_{n}$ are given in (29).
Take arbitrary $\rho>0$ and $1<\sigma^{\prime}<\sigma_{\eta}$. Then we can write $\frac{1}{1-\sigma^{\prime}}=\frac{1}{1-\sigma_{\eta}}+c_{\sigma^{\prime}}$ for suitable $c_{\sigma^{\prime}}>0$. Moreover, we can choose large $|\xi_{0}|$ such that for $e^{-c_{\sigma^{\prime}}W(\ln(1+|\xi|))}<\rho$ when $|\xi|>|\xi_{0}|$. This is due to the fact that $\lim_{\xi\to\infty}e^{-c_{\sigma^{\prime}}W(\ln(1+|\xi|))}=0$.
Therefore, using the property $(W2)$ we obtain
$g_{\sigma_{\eta}}(\xi)$ $=\ln(1+|\xi|)\exp\left\{\frac{1}{\sigma_{\eta}-1}W(\ln(1+|\xi|))\right\}$
$=\ln(1+|\xi|)\exp\left\{\frac{1}{\sigma^{\prime}-1}W(\ln(1+|\xi|))\right\}$
$\,\quad\exp\left\{-c_{\sigma^{\prime}}W(\ln(1+|\xi|))\right\}$
$\leq\rho\ln(1+|\xi|)\exp\left\{\frac{1}{\sigma^{\prime}-1}W(\ln(1+|\xi|))\right\},$
$=\rho g_{\sigma^{\prime}}(\xi),\quad|\xi|>|\xi_{0}|.$ (34)
Choose $n_{3}\in\mathbb{N}$ such that $|\xi_{n}|>|\xi_{0}|$ for $n\geq n_{2}$. Now (30) follows from (33), and (3) for $n^{\prime}_{0}=\max\{n_{1},n_{2}\}$.
The fact that $\varphi\in{{\mathcal{E}}}_{\sigma}\backslash\bigcup_{1<\sigma^{\prime}<\sigma_{\eta}}{{\mathcal{E}}}_{\sigma^{\prime}}$ now follows from Theorem 2.11 (see also Remark 2.12). ∎
Remark 3.4.
Although Theorem 3.3 holds for all $\sigma>1$, the result is optimal when $1<\sigma\leq 2$. This is due to the restriction $\eta>1$ in $\sigma_{\eta}=\frac{\sigma+\eta(\sigma-1)}{1+\eta(\sigma-1)}$, that we needed because constant $C_{\eta}$ that appear in (3) is not finite for $0<\eta\leq 1$.
Finally we can prove Theorem 1.2 as our main result.
Proof of Theorem 1.2.
It is sufficient to construct a low-pass filter $m_{0}\in\mathcal{E}_{\sigma}\setminus\bigcup_{1<\sigma^{\prime}<\sigma}\mathcal{E}_{\sigma^{\prime}}$ that satisfies the assumptions of Theorem 3.3.
Let
$\gamma_{\sigma}(\eta):=\begin{cases}f_{\sigma}(\eta)\quad&,\eta>0\\ 0\quad&,\eta\leq 0,\end{cases}$
where $f_{\sigma}=f_{1,\sigma}$ is given in (3). Note Proposition 2.19 implies that $\gamma_{\sigma}\in\mathcal{E}_{\sigma}(\mathbb{R})$.
Let us define
$\delta_{\sigma}(\xi):=\begin{cases}\left(\int_{0}^{1}\gamma_{\sigma}(\eta)\gamma_{\sigma}(1-\eta)\,d\eta\right)^{-1}\int_{0}^{\xi}\gamma_{\sigma}(\eta)\gamma_{\sigma}(1-\eta)\,d\eta\quad&,\xi>0\\ 0\quad&,\xi\leq 0.\end{cases}$
It is clear that $\delta_{\sigma}(\xi)=1$ for all $\xi\geq 1$, and note $\delta_{\sigma}\in\mathcal{E}_{\sigma}(\mathbb{R})$ (this follows from the property $\widetilde{(M.2)^{\prime}}$ of the $M^{\tau,\sigma}_{p}$). By choosing sufficiently small $\varepsilon>0$ and $0<\rho,\rho^{\prime}<1$, we have
$\gamma_{\sigma}(\eta)$ $\leq f_{\rho,\sigma}(\xi)f_{1-\rho,\sigma}(\eta),$
$\gamma_{\sigma}(1-\eta)$ $\geq f_{\rho^{\prime},\sigma}(\xi)f_{1-\rho^{\prime},\sigma}(1-\eta),\quad\eta\leq\xi<\varepsilon.$ (35)
This follows from the fact that $1-\eta\geq 1-\xi\geq\xi$ when $\xi\in[0,\varepsilon)$, and because $f_{\rho,\sigma}$ are increasing (see part $a)$ of the Lemma (2.16)).
Since $\int_{0}^{\varepsilon}f_{\rho,\sigma}(\eta)d\eta<\infty$, for all $\rho>0$ (see Lemma 2.16 c)), note that (3) implies that there exist $\rho_{1},\rho^{\prime}_{1},C,C^{\prime}>0$ such that for all $\xi\in[0,\varepsilon)$ it holds
$Cf_{\rho_{1},\sigma}(\xi)\leq\delta_{\sigma}(\xi)\leq C^{\prime}f_{\rho^{\prime}_{1},\sigma}(\xi),\quad\xi\in[0,\varepsilon).$
For $\sigma>1$ we define
$\theta_{\sigma}(\xi):=\left(1-\delta_{\sigma}\left(\frac{5\xi-\pi}{3\pi}\right)\right)\delta_{\sigma}\left(\frac{|\xi-\frac{2\pi}{3}|}{d}\right),\quad\xi\in\left[\frac{\pi}{2},\pi\right),\,d\in\left(0,\frac{\pi}{6}\right].$ (36)
This function can be extended on $0,2\pi]$ so that $\theta_{\sigma}(\xi)+\theta_{\sigma}(\xi+\pi)=1$, and then extended further on $\mathbb{R}$ so that it is $2\pi$ periodic.
The desired low-pass filter $m_{0}$ is given by
$m_{0}(\xi)=\sin\left(\frac{\pi}{2}\theta_{\sigma}(\xi)\right),\,\xi\in\mathbb{R}.$ (37)
Indeed, note that Propositions 2.5 (part $iii)$ and part $v)$) and 2.19 imply that $m_{0}\in\mathcal{E}_{\sigma}\setminus\bigcup_{1<\sigma^{\prime}<\sigma}\mathcal{E}_{\sigma^{\prime}}$.
Let us prove that $m_{0}$ in (37) satisfies the assumptions of Theorem 3.3.
For all $\xi\in\mathbb{R}$ it holds
$m_{0}(\xi+\pi)=\sin\left(\frac{\pi}{2}\theta_{\sigma}(\xi+\pi)\right)=\sin\left(\frac{\pi}{2}-\frac{\pi}{2}\theta_{\sigma}(\xi)\right)=\cos\left(\frac{\pi}{2}\theta_{\sigma}(\xi)\right),$
and therefore $m_{0}^{2}(\xi)+m_{0}^{2}(\xi+\pi)=1$, which is condition $iii)$ in (21). Conditions $i)$ and ii)ii) follows directly form the construction.
Note that the functions $m_{0}(\xi)$ and $\theta_{\sigma}(\xi)$ have the same decay rate near $\xi=\frac{2\pi}{3}$ , because $\lim_{\xi\to\frac{2\pi}{3}}\frac{m_{0}(\xi)}{\theta_{\sigma}(\xi)}=\frac{\pi}{2}$. Moreover, $\theta_{\sigma}(\xi)$ satisfies the inequalities (24). This follows since the term $\left(1-\delta_{\sigma}\left(\frac{5\xi-\pi}{3\pi}\right)\right)$ in (36) is bounded on $\mathbb{R}$, and it is strictly positive in the neighborhood of $\xi_{0}=\frac{2\pi}{3}$. Therefore, $m_{0}(\xi)$ also satisfies (24) and we conclude the proof. ∎
For $\sigma>1$ let us define
$N_{x}^{\sigma}:=\left\{f\in L^{2}\left(\mathbb{R}\right)\,\,\Big|\,\,\exp\left\{\rho\,\log^{\frac{\sigma}{\sigma-1}}|\xi|\right\}|\hat{f}(\xi)|\leq C\text{ for some }\rho>0\right\}.$ (38)
In [7, Theorem 4.4] the authors also constructed a wavelet $\psi\in N_{x}^{2}$. To conclude this paper we compare our regularity with the one proposed by $N^{\sigma}_{x}$.
Remark 3.5.
Let $T(k)=\exp\left\{\rho\,\log^{\frac{\sigma}{\sigma-1}}k\right\},\,k>1,\,\sigma>1,$ be the function that describes the decay rate the decay of $\hat{f}$ in (38). Then by the calculation done in [13, Section 4.1], it follows that $T(k)$ is associated function for $M_{p}=q^{p^{\sigma}}$, for suitable $q>1$ when $1<\sigma\leq 2$.
In particular we can prove the following Lemma.
Lemma 3.6.
Let $N_{x}^{\sigma}$ be defined in (38), and
$\mathcal{E}_{\sigma}^{\Gamma}=\{f\in L^{2}(\mathbb{R})\mid\hat{f}\in{{\mathbf{\Gamma}}}_{\sigma}(\mathbb{R})\},$
for every $\sigma>1$, where ${\mathbf{\Gamma}}_{\sigma}(\mathbb{R})$ is given in the Definition 2.10. Then
$N_{x}^{\sigma}\subset\mathcal{E}_{\sigma}^{\Gamma}\subset\mathcal{E}_{\sigma}.$
Proof.
Let $\sigma>1$ and $|\xi_{0}|>\frac{1+\sqrt{5}}{2}$ be fixed. Then for $|\xi|>|\xi_{0}|$ we have
$\log^{\frac{\sigma}{\sigma-1}}|\xi|$ $=W^{\frac{1}{\sigma-1}}(\log|\xi_{0}|)\frac{1}{W^{\frac{1}{\sigma-1}}(\log|\xi_{0}|)}\log^{\frac{\sigma}{\sigma-1}}|\xi|$
$\geq\frac{W^{\frac{1}{\sigma-1}}(\log|\xi_{0}|)}{2^{\frac{\sigma}{\sigma-1}}}\frac{\log^{\frac{\sigma}{\sigma-1}}(|\xi|^{2})}{W^{\frac{1}{\sigma-1}}(\log(1+|\xi|))}$
$\geq\frac{W^{\frac{1}{\sigma-1}}(\log|\xi_{0}|)}{2^{\frac{\sigma}{\sigma-1}}}\frac{\log^{\frac{\sigma}{\sigma-1}}(1+|\xi|)}{W^{\frac{1}{\sigma-1}}(\log(1+|\xi|))},$
and therefore, if $f\in N_{x}^{\sigma}$ we conclude
$|\hat{f}(\xi)|\leq Ce^{-\rho\,\log^{\frac{\sigma}{\sigma-1}}|\xi|}\leq Ce^{-\rho^{\prime}g_{\sigma}(\xi)},$
for suitable $C,\rho^{\prime}>0$ and $|\xi|$ large enough. This proves that $f\in\mathcal{E}_{\sigma}^{\Gamma}$. Inclusion $\mathcal{E}_{\sigma}^{\Gamma}\subset\mathcal{E}_{\sigma}$ follows directly from the Theorem 2.11 b). ∎
3.1 Illustrations
In this section, we present illustrations of the low-pass filter constructed in the proof of Theorem 1.2, together with the corresponding scaling function and wavelet. All figures were generated in Matlab. The parameters used in (36) and (37) are $\sigma=2$ and $d=\frac{\pi}{12}$.
Figure 1: The plot of the low-pass filter $m_{0}$ given in (37). Note that it features a small bump near $\tfrac{2\pi}{3}$, which is a consequence of controlling its local decay.
Figure 2: Plot of the Fourier transform of the scaling function; $\hat{\varphi}$.
Figure 3: Note that $\hat{\varphi}$ inherits the bumps of $m_{0}$ near the invariant cycle points. Therefore $\hat{\varphi}$ is not band-limited.
Figure 4: Plot of Fourier transform of the wavelet; $\hat{\psi}$
Figure 5: Plot of the wavelet $\psi$
4 Appendix
Proof of Lemma 2.17.
We prove that $\lim_{x\to 0^{+}}f_{\rho_{\sigma},\sigma}^{(j)}\left(x\right)=0$, where $f_{\rho_{\sigma},\sigma}$ is given in (18).
For $x>0$ and $j\in\mathbb{N}$ we have
$f_{\rho_{\sigma},\sigma}^{(j)}\left(x\right)=\left(e^{-\rho_{\sigma}g_{\sigma}\left(\frac{1}{x}\right)}\right)^{(j)}=\left(\left(e^{-\rho_{\sigma}\,\cdot\,}\circ\omega_{\sigma}(\,\cdot\,)\circ\ln\left(1+\frac{1}{\cdot}\right)\right)(x)\right)^{(j)},$ (39)
where $g_{\sigma}$ and $\omega_{\sigma}$ are given in (3).
Note that
$\left(\ln\left(1+\frac{1}{x}\right)\right)^{(n)}=(-1)^{n-1}(n-1)!\left(\frac{1}{(x+1)^{n}}-\frac{1}{x^{n}}\right),\quad x>0,\;n\in\mathbb{N}.$ (40)
Let us compute the $n$-th derivative of $\omega_{\sigma}(x)$. Using Leibniz product formula we obtain
$\omega_{\sigma}^{(n)}(x)$ $=\sum_{k=0}^{n}\binom{n}{k}\left(x^{\frac{\sigma}{\sigma-1}}\right)^{(n-k)}\left(\frac{1}{W^{\frac{1}{\sigma-1}}(x)}\right)^{(k)}$
$=\sum_{k=0}^{n}\binom{n}{k}\prod_{l=0}^{n-k-1}\left(\frac{\sigma}{\sigma-1}-l\right)x^{\frac{\sigma}{\sigma-1}-n+k}\left(\frac{1}{W^{\frac{1}{\sigma-1}}(x)}\right)^{(k)},\,x>0,\,n\in\mathbb{N}.$
By Faá di Bruno formula and (W4), for $0\leq k\leq n$ we have
$\left(\frac{1}{W^{\frac{1}{\sigma-1}}(x)}\right)^{(k)}=\sum_{\begin{subarray}{c}m_{1},\dots,m_{k}\in\mathbb{N}\\ m_{1}+2m_{2}+\dots+km_{k}=k\end{subarray}}\frac{k!(-1)^{m_{1}+m_{2}+\dots+m_{k}}}{m_{1}!m_{2}!\dots m_{k}!}\times$
$\times\prod_{j=0}^{m_{1}+\dots+m_{k}-1}\left(\frac{1}{\sigma-1}+j\right)\frac{1}{W(x)^{\frac{1}{\sigma-1}+m_{1}+\dots+m_{k}}}\prod_{s=1}^{k}\frac{1}{s!^{m_{s}}}\left(\frac{W^{s}(x)p_{s}(W(x))}{x^{s}(1+W(x))^{2s-1}}\right)^{m_{s}}$
$=\frac{1}{x^{k}}\frac{W(x)^{k-\frac{1}{\sigma-1}}}{(1+W(x))^{2k}}\sum_{\begin{subarray}{c}m_{1},\dots,m_{k}\in\mathbb{N}\\ m_{1}+2m_{2}+\dots+km_{k}=k\end{subarray}}A_{k}(m_{1},\dots,m_{k})\left(\frac{1+W(x)}{W(x)}\right)^{m_{1}+\dots+m_{k}}\times$
$\times\prod_{s=1}^{k}(p_{s}(W(x)))^{m_{s}},\quad x>0,$
where
$A_{k}(m_{1},\dots,m_{k})=\frac{(-1)^{m_{1}+m_{2}+\dots+m_{k}}k!}{m_{1}!m_{2}!2!^{m_{2}}\dots m_{k}!k!^{m_{k}}}\prod_{j=0}^{m_{1}+\dots+m_{k}-1}\left(\frac{1}{\sigma-1}+j\right).$
Therefore,
$\omega_{\sigma}^{(n)}(x)=\sum_{k=0}^{n}a_{k,n}(\sigma)\,x^{\frac{\sigma}{\sigma-1}-n+k}\frac{1}{x^{k}}\frac{W^{k-\frac{1}{\sigma-1}}(x)}{(1+W(x))^{2k}}$
$\quad\sum_{\begin{subarray}{c}m_{1},\dots,m_{k}\in\mathbb{N}\\ m_{1}+2m_{2}+\dots+km_{k}=k\end{subarray}}A_{k}(m_{1},\dots,m_{k})\left(\frac{1+W(x)}{W(x)}\right)^{m_{1}+\dots+m_{k}}\prod_{s=1}^{k}(p_{s}(W(x)))^{m_{s}}$
$=x^{\frac{\sigma}{\sigma-1}-n}\sum_{k=0}^{n}\sum_{\begin{subarray}{c}m_{1},\dots,m_{k}\in\mathbb{N}\\ m_{1}+2m_{2}+\dots+km_{k}=k\end{subarray}}a_{k,n}(\sigma)A_{k}(m_{1},\dots,m_{k})\,\frac{W^{k-\frac{1}{\sigma-1}}(x)}{(1+W(x))^{2k}}\times$
$\quad\times\left(\frac{1+W(x)}{W(x)}\right)^{m_{1}+\dots+m_{k}}\prod_{s=1}^{k}(p_{s}(W(x)))^{m_{s}},\quad x>0,\;n\in\mathbb{N},$ (41)
where $a_{k,n}(\sigma)=\binom{n}{k}\prod_{l=0}^{n-k-1}\left(\frac{\sigma}{\sigma-1}-l\right)$.
To avoid further technicalities, we observe that (40), (4), and successive application of the Faá di Bruno formula, imply that (39) can be written as (finite) sum of terms
$\alpha_{\sigma}e^{-\rho_{\sigma}g_{\sigma}\left(\frac{1}{x}\right)}\left(\ln\left(1+\frac{1}{x}\right)\right)^{a_{\sigma}}\frac{W^{b_{\sigma}}\left(\ln\left(1+\frac{1}{x}\right)\right)}{\left(1+W\left(\ln\left(1+\frac{1}{x}\right)\right)\right)^{c}}\frac{1}{x^{d^{\prime}}(x+1)^{d^{\prime\prime}}},\quad x>0,$ (42)
for suitable $\alpha_{\sigma},a_{\sigma},b_{\sigma}\in\mathbb{R},\,c,d^{\prime},d^{\prime\prime}\geq 0$, where we also recall that $p_{s}(W(x))$ appearing in (4) is polynomial with respect to $W(x)$.
Setting $s=\ln\left(1+\frac{1}{x}\right)$ in (42) and using (5), we obtain
$e^{-\rho_{\sigma}\omega_{\sigma}(s)}s^{a_{\sigma}}\frac{W^{b_{\sigma}}(s)}{(1+W(s))^{c}}\frac{(e^{s}-1)^{d^{\prime}+d^{\prime\prime}}}{(e^{s})^{d^{\prime\prime}}}\sim e^{-\rho_{\sigma}\omega_{\sigma}(s)}s^{a_{\sigma}}\ln^{b_{\sigma}-c}(s)e^{d^{\prime}s},$ (43)
when $s\to\infty$.
Note that there exist constants $\beta,d_{0}>0$ such that
$|s^{a_{\sigma}}\ln^{b_{\sigma}-c}(s)e^{d^{\prime}s}|\leq\beta e^{d_{0}s},\quad s>0,$
hence part $b)$ of the Lemma 2.16 and (43) imply that each term of the form (42) tends to 0 when $s\to+\infty$. Finally, this proves $\lim_{x\to 0^{+}}f_{\rho_{\sigma},\sigma}^{(j)}(x)=0$. ∎
5 Acknowledgments
This research was supported by the Science Fund of the Republic of Serbia, #\#GRANT No. 2727, Global and local analysis of operators and distributions - GOALS. The authors were also supported by the Ministry of Science, Technological Development, and Innovation of the Republic of Serbia - the second and third authors by Grants No. 451-03-137/2025-03/200125 and 451-03-136/2025-03/200125, and the first author by Grant No. 451-03-136/2025-03/200156.